a 185 lb jogger is running at 8 mph. as he nears the end of his run he decides to sprint increasing his speed to 18 mph. how does his kinetic energy change?
Answer #1

116 = 0.454 kg mass me 185 W = 185*0.454 kg = 83.99kg Vi = 8 mph = 8x0.447 = 3.576 ms z 18 mph = 1840.447 - 8.046 m/s ARE - m